OpenStudy (anonymous):

Solve for x: −4|x + 5| = −16

OpenStudy (lynfran):

-4|x+5|=-16 |x+5|=-16/-4 |x+5|=4 .... can u solve for x?

OpenStudy (lynfran):

x+5=-4 and x+5=4 so....x=-4-5 and x=4-5 so whats x=?

OpenStudy (lynfran):

@bestleah

OpenStudy (lynfran):

whats -4-5=? and whats 4-5=?

OpenStudy (lynfran):

ok x=-9 and x=-1 are ur solutions ok

OpenStudy (anonymous):

Thanks